OpenStudy (anonymous):

*PLEASE HELP* Identify the slope and the y-intercept of 4x+8y=8?

OpenStudy (anonymous):

you should just have to solve for y

OpenStudy (anonymous):

y=-1/2x + 1..I believe. Check to make sure

OpenStudy (anonymous):

slope is -1/2 y-int. is 1

OpenStudy (anonymous):

4x + 8y = 8 (subtract 4x from both sides) 8y = -4x + 8 (now divide by 8) y = -1/2x + 1 In y = mx + b form, the m position is the slope and the b position is the y intercept. You are correct brianhotz ... slope = -1/2 and y int = 1